OpenStudy (anonymous):

Rick and James live 285 miles apart. At the same time, they start driving toward each other on the same road. Rick’s constant rate is 40 mph. James’ is 55 mph. How long will it take them to meet?

OpenStudy (studygurl14):

just google it

OpenStudy (studygurl14):

also, make sure to doubel check any answer you get from the internet, cause they might be bad

OpenStudy (jonnyvonny):

40+55=95 285/95=??

OpenStudy (cwrw238):

the speed of approach is 40 + 55 = 95 mph

OpenStudy (anonymous):

thankies

OpenStudy (cwrw238):

usethe formula time = distance / speed

OpenStudy (anonymous):

you should study!